OpenStudy (anonymous):

Write a recursive formula for the sequence 7, 13, 19, 25, 31, ... Then find the next term.

OpenStudy (walters):

6n+1 next term is 37

OpenStudy (anonymous):

\[a_1 = 7\]\[a_2 = a_1 + 6\]\[a_3 = a_2 + 6\]the recursive formula:\[a_n = a_{n - 1} + 6\]the next term\[a_6 = a_5 + 6\]\[a_6 = 31 + 6\]\[a_6 = 37\]